# The temperature and metallicity distributions of the ICM: insights with TNG-Cluster for XRISM-like observations

### v1

@@ Acknowledgements
DC is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD) and, together with AP, acknowledges funding from the European Union (ERC, COSMIC-KEY, 101087822, PI: Pillepich). DN acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) through an Emmy Noether Research Group (grant number NE 2441/1-1).
@@
Appendix A Spectral emission model systematic uncertainties
This section reports in detail the systematic uncertainties, associated with the choice of spectral emission model, on the inference of the true properties of the ICM, as predicted by TNG-Cluster.
Inferred - True value | bvapec | double bvapec | bvlognorm | bvgadem |
@@ Appendix A Spectral emission model systematic uncertainties
Mass-weighted central gas (all) | -1.05 (-0.75,+0.74) | -1.26 (-0.89,+1.17) | -1.21 (-0.73,+0.72) | -1.22 (-0.74,+0.77) |
Mass-weighted central gas (accepted only) | -1.05 (-0.75,+0.74) | -0.64 (-0.89,+0.83) | -1.14 (-0.75,+0.63) | -1.06 (-0.66,+0.63) |
Emission-weighted FoV gas (all) | 0.09 (-0.24,+0.16) | -0.12 (-0.52,+0.81) | -0.08 (-0.21,+0.23) | -0.09 (-0.17,+0.24) |
Emission-weighted FoV (accepted only) | 0.09 (-0.24,+0.16) | 0.23 (-0.89,+0.70) | -0.14 (-0.18,+0.13) | -0.06 (-0.16,+0.14) |
Thermal Width Residuals () (2) | ||||
Mass-weighted central gas (all) | 0.40 (-0.67,+1.80) | 0.92 (-1.04,+1.50) | ||
Mass-weighted central gas (accepted only) | 0.63 (-1.18,+2.23) | 1.51 (-1.92,+1.74) | ||
Emission-weighted FoV gas (all) | -1.39 (-0.01,+2.41) | 0.84 (-1.01,+1.24) | ||
Emission-weighted FoV gas (accepted only) | -2.12 (-0.01,+3.59) | 1.46 (-1.55,+1.62) | ||
@@
Mass-weighted central gas (all) | -0.13 (-0.11,+0.10) | -0.12 (-0.11,+0.09) | -0.18 (-0.11,+0.09) | -0.12 (-0.11,+0.10) |
Mass-weighted central gas (accepted only) | -0.13(-0.11,+0.10) | -0.08 (-0.05,+0.07) | -0.15 (-0.05,+0.07) | -0.08 (-0.04,+0.06) |
Emission-weighted FoV gas (all) | -0.02 (-0.02,+0.03) | -0.01 (-0.02,+0.03) | -0.06 (-0.02,+0.03) | -0.01 (-0.02,+0.03) |
Emission-weighted FoV gas (accepted only) | -0.02 (-0.02,+0.03) | -0.00 (-0.02,+0.03) | -0.06 (-0.02,+0.03) | -0.00 (-0.02,+0.03) |
Inferred / True value | bvapec | double bvapec | bvlognorm | bvgadem |
@@
Mass-weighted central gas (all) | 0.86 (-0.08,+0.06) | 0.82 (-0.12,+0.16) | 0.82 (-0.08,+0.07) | 0.82 (-0.07,+0.09) |
Mass-weighted central gas (accepted only) | 0.86 (-0.08,+0.06) | 0.88 (-0.20,+0.17) | 0.82 (-0.09,+0.08) | 0.83 (-0.08,+0.07) |
Emission-weighted FoV gas (all) | 1.05 (-0.07,+0.02) | 0.99 (-0.13,+0.18) | 0.99 (-0.05,+0.04) | 0.99 (-0.03,+0.04) |
Emission-weighted FoV gas (accepted only) | 1.05 (-0.07,+0.02) | 1.06 (-0.24,+0.17) | 0.97 (-0.04,+0.03) | 0.99 (-0.03,+0.03) |
@@
Mass-weighted central gas (all) | 0.75 (-0.19,+0.16) | 0.78 (-0.20,+0.16) | 0.64 (-0.17,+0.15) | 0.77 (-0.20,+0.16) |
Mass-weighted central gas (accepted only) | 0.75 (-0.19,+0.16) | 0.85 (-0.14,+0.12) | 0.70 (-0.12,+0.10) | 0.86 (-0.12,+0.10) |
Emission-weighted FoV gas (all) | 0.96 (-0.08,+0.08) | 1.00 (-0.09,+0.08) | 0.82 (-0.08,+0.07) | 1.00 (-0.08,+0.09) |
Emission-weighted FoV gas (accepted only) | 0.96 (-0.08,+0.08) | 1.02 (-0.06,-0.08) | 0.85 (-0.05,+0.06) | 1.03 (-0.06,+0.06) |

### v2

@@ Acknowledgements
We thank the reviewer for their constructive and useful suggestions. DC and AP thank Joey Braspenning for useful comments on the paper. DC is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD) and, together with AP, acknowledges funding from the European Union (ERC, COSMIC-KEY, 101087822, PI: Pillepich). DN acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) through an Emmy Noether Research Group (grant number NE 2441/1-1).
@@
Appendix A spectral-emission model systematic uncertainties
This section reports in detail, in a series of Tables (2, 3), the systematic uncertainties, associated with the choice of spectral-emission model, on the inference of the true properties of the ICM, as predicted by TNG-Cluster.
| Inferred - True value | bvapec | double bvapec | bvgadem | bvlognorm |
@@
| Mass-weighted central gas (all) | -1.05 (-0.75,+0.74) | -1.32 (-1.07,+1.22) | -1.18 (-0.78,+0.74) | -1.21 (-0.73,+0.72) |
| Mass-weighted central gas (accepted only) | -1.09 (-0.72,+0.72) | -0.68 (-0.92,+0.87) | -0.98 (-0.59,+0.54) | -1.17 (-0.67,+0.68) |
| emission-weighted FoV gas (all) | 0.09 (-0.24,+0.16) | -0.17 (-0.58,+0.86) | -0.05 (-0.17,+0.22) | -0.08 (-0.22,+0.24) |
| emission-weighted FoV (accepted only) | 0.06 (-0.21,+0.15) | 0.22 (-0.95,+0.69) | 0.00 (-0.16,+0.11) | -0.15 (-0.17,+0.14) |
| Temperature width Residuals () (2) | ||||
| Mass-weighted central gas (all) | 1.56 (-1.77,+0.55) | 1.05 (-1.28,+0.70) | ||
| Mass-weighted central gas (accepted only) | 2.08 (-1.73,+0.64) | 1.51 (-1.21,+0.79) | ||
| emission-weighted FoV gas (all) | 1.27 (-1.67,+0.56) | 0.41 (-1.05,+0.85) | ||
| emission-weighted FoV gas (accepted only) | 1.83 (-1.58,+0.81) | 0.81 (-0.89,+0.92) | ||
@@
| Mass-weighted central gas (all) | -0.13 (-0.11,+0.10) | -0.12 (-0.11,+0.09) | -0.12 (-0.12,+0.09) | -0.12 (-0.12,+0.09) |
| Mass-weighted central gas (accepted only) | -0.14 (-0.11,+0.10) | -0.09 (-0.05,+0.07) | -0.08 (-0.04,+0.06) | -0.08 (-0.04,+0.07) |
| emission-weighted FoV gas (all) | -0.02 (-0.02,+0.03) | -0.01 (-0.02,+0.03) | -0.01 (-0.02,+0.03) | -0.00 (-0.02,+0.03) |
| emission-weighted FoV gas (accepted only) | -0.02 (-0.02,+0.03) | 0.01 (-0.01,+0.03) | 0.01 (-0.01,+0.03) | -0.01 (-0.01,+0.03) |
| Inferred / True value | bvapec | double bvapec | bvgadem | bvlognorm |
@@
| Mass-weighted central gas (all) | -0.14 (-0.08,+0.06) | -0.19 (-0.12,+0.17) | -0.18 (-0.06,+0.09) | -0.18 (-0.08,+0.07) |
| Mass-weighted central gas (accepted only) | -0.15 (-0.07,+0.06) | -0.11 (-0.18,+0.16) | -0.15 (-0.06,+0.06) | -0.19 (-0.08,+0.08) |
| emission-weighted FoV gas (all) | 0.05 (-0.07,+0.02) | -0.02 (-0.14,+0.20) | -0.01 (-0.03,+0.04) | -0.01 (-0.05,+0.04) |
| emission-weighted FoV gas (accepted only) | 0.04 (-0.06,+0.01) | 0.06 (-0.23,+0.17) | 0.00 (-0.03,+0.02) | -0.03 (-0.04,+0.03) |
@@
| Mass-weighted central gas (all) | -0.25 (-0.19,+0.17) | -0.22 (-0.20,+0.20) | -0.23 (-0.20,+0.20) | -0.22 (-0.20,+0.17) |
| Mass-weighted central gas (accepted only) | -0.26 (-0.19,+0.16) | -0.16 (-0.13,+0.13) | -0.15 (-0.11,+0.10) | -0.15 (-0.12,+0.11) |
| emission-weighted FoV gas (all) | -0.04 (-0.08,+0.04) | 0.00 (-0.09,+0.09) | -0.01 (-0.09,+0.09) | 0.01 (-0.08,+0.10) |
| emission-weighted FoV gas (accepted only) | -0.03 (-0.07,+0.07) | 0.03 (-0.05,+0.07) | 0.03 (-0.04,+0.06) | 0.05 (-0.06,+0.07) |
Appendix B Inferred LoS velocity dispersion
In this section, we report, in addition to the temperature and metallicity properties of the ICM – the focus of the paper –, also the LoS velocity dispersion values inferred according to each spectral-emission model. We do so for the full sample of studied clusters. As previously mentioned, comparing these values to the intrinsic 3D velocity dispersion of the gas cells from the simulation exceeds the scope of this work. However, considering potential biases due to degeneracies among all free parameters and considering the current interest in the kinematics of cluster centres enables by XRISM, we want to provide results also on our kinematic inferences.
In Figure 15 we present the inferred velocity dispersions for all spectral-emission models used in this work as a function of redshift. We note that we find little to no difference in the inferred values between models, with the possible exception of the double bvapec model for low mass halos. This suggests that the inferred kinematics appear to be, for the most part or at least for the average cluster, unaffected by the parametrisation of the ICM’s thermal distribution.
For comparison, we also include in Figure 15 the observationally-inferred velocity dispersions from the cores of a few selected galaxy clusters, as available from the current literature (large black markers, from lower to higher masses: Centaurus, Hydra A, Perseus, Abell 2029, Coma and Ophiuchus). We see that observations are consistent with the TNG-Cluster predictions, deviating by less than 1 from the average trends. The only outlier appears to be Ophiuchus, at the highest mass end. However, in the observational setup, the central region only covers the central 25 kpc, compared to the 33.5 kpc used in this work, and the velocity dispersion appears to increase with radius (Fujita et al., 2025). This might imply that the velocity dispersion of Ophiuchus may also fall within 1 of the TNG-Cluster predictions, if the central pointings were matched. This preliminary and at-face-value comparison may serve as a useful starting point for future, more tailored analyses (e.g. Xrism Collaboration et al., 2025b, c).
Appendix C On fitting the temperature distributions of simulated ICM
As discussed in Section 3.1, we fit the intrinsic temperature distribution of the ICM predicted by TNG-Cluster with both a Normal and a Log-Normal distribution. In Figure 16 we present the residuals, i.e. the difference between intrinsic and modelled probability mass functions, for both fit results. On average (thick curves), the offsets do not exceed the 1 per-cent level, suggesting that both Normal and a Log-Normal functional forms can provide a reasonable description of the thermal properties of the gas across our entire cluster sample.
However, the tensions are more significant for individual halos (thin curves), with a large cluster-to-cluster variation reflected in the large scatter of values. As we have already highlighted for the entire sample, the tensions primarily arise from highly asymmetric intrinsic distributions, due to extended tails of either hotter or cooler gas, relative to the mode of the distribution. We note that for both functional forms, we find an increase in the residuals on the sides of the temperature peak due to over-/under-fitting, as explained in the text.